The total height of a building and the flagpole on the roof is 147 feet. The building is 6 times as high as the flagpole. How high is the flagpole?
Answer by mananth(16949)

The total height of a building and the flagpole on the roof is 147 feet. The building is 6 times as high as the flagpole. How high is the flagpole?
...
flagpole = x feet
building = 147-x
..
147-x= 6x
add x to both sides
147= 6x+x7x=147
/7
x= 21 feet the height of the flagpole
...
m.ananth@hotmail.ca